### v2

# Summarization of Opinionated Political Documents with Varied Perspectives

Abstract
Global partisan hostility and polarization has increased, and this polarization is heightened around presidential elections. Models capable of generating accurate summaries of diverse perspectives can help reduce such polarization by exposing users to alternative perspectives. In this work, we introduce a novel dataset and task for independently summarizing each political perspective in a set of passages from opinionated news articles. For this task, we propose a framework for evaluating different dimensions of perspective summary performance. We benchmark 11 summarization models and LLMs of varying sizes and architectures through both automatic and human evaluation. While recent models like GPT-4o perform well on this task, we find that all models struggle to generate summaries that are faithful to the intended perspective. Our analysis of summaries focuses on how extraction behavior is impacted by features of the input documents [^1] We make our code available at https://github.com/NickDeas/PoliSum.
Summarization of Opinionated Political Documents with Varied Perspectives
Nicholas Deas, Kathleen McKeown Department of Computer Science, Columbia University [ndeas,kathy]@cs.columbia.edu
1 Introduction
Political ideologies can lead people to develop misperceptions of groups with opposing opinions Chambers et al. (2006), and political events, such as the 2024 US presidential election, French legislative election, or the Brexit referendum, can reinforce negative attitudes Hanna et al. (2013); Wellings et al. (2024). These misperceptions, reinforced by news consumption on social media Levy (2021), contribute to increased polarization and instability Braley et al. (2023); Lees and Cikara (2021). Exposure to alternative perspectives, however, has been shown to help alleviate polarization Balietti et al. (2021). To encourage such exposure, some groups focus on aggregating (e.g., All Sides [^2] https://www.allsides.com/) or summarizing different political perspectives on divisive issues (e.g., The Flip Side [^3] https://www.TheFlipSide.io/, Ground News [^4] https://ground.news/).
Large language models’ (LLMs) ability to summarize opinions and news has recently neared human performance Zhang et al. (2023a); Bhaskar et al. (2023). Recent work, however, has also shown that LLMs can unfairly represent diverse opinions in review and tweet summarization Zhang et al. (2023b); Huang et al. (2023); Tay (2019).
Figure 1: Example of the Mixed-Perspective Setting of PoliSum. Given a set of opinionated passages reflecting a mix of views, PoliSum requires models to generate both a summary of the left and right-leaning political perspectives.
To evaluate current models’ capabilities in summarizing multiple perspectives in a set of input documents, we propose a novel dataset, PoliSum, and task, mixed perspective summarization [^5] We make our dataset available to researchers who have been granted permission to use the data for their work by The Flip Side.. Our task involves generating independent abstractive summaries of both left and right-leaning political perspectives given collections of opinionated news passages with mixed perspectives. Mixed perspective summarization thus requires models to both distinguish the perspectives reflected in input texts as well as produce a pair of summaries representing those differing perspectives. We summarize our primary contributions as follows:
  1. 1.
    We introduce a novel benchmark dataset, PoliSum, for mixed perspective summarization, requiring models to independently summarize political perspectives on controversial issues, provided passages from news and op-eds.
  2. 2.
    We introduce an initial framework for evaluating different components of perspective summaries, including both human evaluation and automatic metrics.
  3. 3.
    We benchmark 11 models, finding that most approaches struggle to faithfully summarize the intended perspective and that improving automatic evaluation metrics is an important area for future work.
  4. 4.
    We analyze model extraction behavior (e.g., which documents models extract from and to what degree) and find that generated summaries suffer from biases due to input document position, document length, and use of arousing terms.
2 PoliSum Dataset
Dataset Domain Multi-document Multi-summary # Gold Gold Length % Novel n-grams
1 2 3
XSUM News 11,334 23.3 35.8 83.5 95.5
Multi-News News 5,622 263.7 17.8 57.1 75.7
CoCoTrip Reviews 48348$\cdot$ 348 ⋅ 3 132.9 22.8 72.4 91.4
PoliSum Op-Eds 9072907$\cdot$ 2907 ⋅ 2 19.7 40.4 86.8 97.5
Table 1: Summary statistics for PoliSum and existing summarization datasets, including data size, reference length, and percentage of novel n-grams in the summary compared to source documents. # Gold for multi-summary datasets is presented as # Samples $\cdot$ # References/Sample. CoCoTrip statistics referenced from Iso et al. (2022).
Data Source. To enable evaluating political perspective summaries, we collect opinionated texts and paired summaries from The Flip Side [^6] We thank The Flip Side for generously allowing us to use their archive for this work: https://www.TheFlipSide.io/archives. Since September 2017, The Flip Side has covered politically controversial issues (i.e. overturning of Roe v. Wade, presidential elections) in the United States. Toward bridging political divides, each publication includes a headline reflecting the central issue, curated passages authored by political commentators with varying political leaning, and a pair of short summaries of the left and right perspectives [^7] Some publications lack individual perspective summaries, primarily on topics where both parties largely agree or for meta-articles that do not cover political topics. These samples are not included in the dataset.. We collect this information from The Flip Side archive, gathering nearly 1,000 samples. An excerpted example of source passages and summaries is shown in Table 2.
Headline Hollywood Strike
Perspective Left Right
Source Passages …This is a huge issue, but there’s another one — AI… A lot of the TV episodes and movies produced by Hollywood are, by nature, highly formulaic… …By replacing the once majority straight white cast with a ridiculously disproportionate percentage of minorities, industry players alienated a commensurate portion of the audience…
Summaries The left supports the strikers, arguing that AI will soon be a threat to workers in many industries. The right argues that Hollywood’s troubles stem from its embrace of leftist politics.
Table 2: Example of partial input passages and paired perspective summaries in PoliSum.
Table 1 presents summary statistics of the collected multi-document, multi-summary dataset (data statement included in Appendix C). We also show statistics for CoCoTrip (Iso et al. 2022; the most comparable multi-document, multi-summary dataset), XSUM (Narayan et al. 2018; a single-document, single-summary corpus with similarly short references), and Multi-News (Fabbri et al. 2019, a multi-document, single-summary news corpus). The aim of CoCoTrip is to generate multiple hotel review summaries for a pair of comparable hotels: two summaries highlighting each hotel’s unique attributes, and one highlighting similarities. In contrast, PoliSum contains almost 20 times more labeled examples and focuses on political perspectives rather than reviews. Additionally, compared to XSum, PoliSum is similar in length but slightly more abstractive, measured by n-gram overlap between source and reference. On average, there are 9.6 source passages for each publication (4.8 per perspective), with a combined length comparable to a news article ($\sim$1,300 tokens).
Perspective Summarization Task. Using this dataset, the aim of the our proposed task is to generate a pair of summaries (i.e., Summaries in Table 2) representing the left and right political perspectives on a controversial topic (i.e. Headline). As input, models are provided with two sets of opinionated editorial passages (i.e., Source Passages), each representing one of the two political perspectives. Thus, the dataset enables a multi-document, multi-summary task. As with traditional news summarization tasks, the summary for each perspective should condense the information present in the passages. At the same time, the generated left and right-perspective summaries should faithfully represent the perspectives reflected in their respective passages.
3 Experiments
Traditional opinion summarization typically involves summarizing a common perspective among all inputs, which we refer to as the Single-Perspective setting [^8] Experimental details and results of the Single-Perspective task on PoliSum are included in Appendix E. We instead introduce an additional challenge requiring models to identify input documents relevant to the intended perspective, referred to as the Mixed-Perspective setting and illustrated in Figure 1. In this setting, approaches are provided with a mixed set of left and right-leaning passages as input and are tasked with producing independent left and right-perspective summaries from the same input. On average, sets of inputs contain $\sim$10 passages total, $\sim$5 passages for each perspective.
Model inputs are constructed by concatenating all passages, ensuring that the left and right perspectives alternate when possible. Due to position biases in summarization Chhabra et al. (2024), the perspective reflected in the first input document may spuriously impact the resulting summary quality. To mitigate these effects, we report average performance between two variations of the input data. First, models are evaluated on inputs beginning with a left-perspective passage followed by alternating perspectives (e.g., left, right, left, etc.), and then on inputs beginning with a right-perspective passage followed by alternating perspectives (e.g., right, left, right, etc.). The order of individual passages that share a perspective is randomized. Because the reference summaries are limited to a single sentence, all model generations are truncated to the first sentence before evaluation.
3.1 Models
A variety of summarization models and instruction-tuned LLMs are considered.
Extractive Upper Bound. To represent potential extractive summarization approaches, we estimate an upper bound performance based on semantic similarity to the reference summaries. We consider all sentences contained in the source passages to be potential extractive summaries. We then select the sentence with the highest BERTScore relative to the reference summary and calculate the remaining coverage and faithfulness metrics.
Neural Summarization Baselines. To represent neural summarization models, we evaluate the large variants of BART Lewis et al. (2019), T5 Raffel et al. (2019), and Flan-T5 Chung et al. (2022). As BART is not pretrained on any summarization task, we use BART finetuned on the CNN/DailyMail (CNN/DM; Hermann et al. 2015; Nallapati et al. 2016) dataset. In contrast to PoliSum, CNN/DM and other prior summarization datasets lack summaries for individual perspectives. So, in order to encourage BART to generate distinct summaries matching the references’ format, the BART decoder is forced to begin generations with "The left" or "The right" respectively. T5 and Flan-T5 are prompted to generate summaries of individual perspectives. See Appendix A for other generation hyperparameter details.
LLMs. We evaluate 8 additional LLMs of varying sizes and architectures to establish baselines for current state-of-the-art models [^9] We use instruction-tuned variants of all applicable LLMs, but for simplicity, omit ”Instruct” in names throughout results.. Mistral-7B (mistralai/Mistral-7B-Instruct-v0.1; Jiang et al. 2023), Llama-3.0-8B and 70B (meta-llama/Meta-Llama-3-8B-Instruct and meta-llama/Meta-Llama-3-70B-Instruct; Dubey et al. 2024), Llama-3.3-70B (meta-llama/Meta-Llama-3.3-70B-Instruct), Vicuna-7B and 13B (lmsys/vicuna-7b-v1.5 and lmsys/vicuna-13b-v1.5; Chiang et al. 2023), Mixtral-8x22B (mistralai/Mixtral-8x22B-Instruct-v0.1; Jiang et al. 2024), and GPT-4o (gpt-4o-2024-05-13) [^10] https://openai.com/index/hello-gpt-4o/. Due to the length of input documents, we evaluate all LLMs in a zero-shot setting. See Appendix B for additional generation hyperparameters, prompts, and reproducibility details.
3.2 Metrics
Coverage.
Summary coverage is first measured with Rouge-1 and Rouge-L between the generated perspective summaries and their corresponding references Lin (2004a). To account for the weaknesses of Rouge, we also employ BERTScore Zhang et al. (2020) and Bleurt Sellam et al. (2020) using the deberta-large-xnli and Bleurt-20-D6 checkpoints respectively due to higher correlations with human judgments compared to alternatives [^11] See human correlations for BERTScore linked in https://github.com/Tiiiger/bert_score and for Bleurt in https://github.com/google-research/Bleurt/blob/master/checkpoints.md. For each metric, we also use a two-sample t-test to determine whether left and right-perspective coverage scores significantly differ ($p\leq 0.05$).
Faithfulness.
In addition to coverage, we also measure how well the generated summaries are consistent and faithful to the input passages representing the intended perspective. Because faithfulness metrics for our specific task do not exist, we rely on existing factuality metrics as approximations: SummaC Laban et al. (2022) and AlignScore Zha et al. (2023). SummaC is a NLI-based metric that measures consistency between summaries and source documents. Alternatively, AlignScore uses information alignment to measure factual consistency. All metrics produce a score between 0 and 100 where higher values reflect more consistent and faithful summaries.
Perspective.
Figure 2: Example perspective summary with indicators of different perspective dimensions highlighted: Consensus, Stance and Intensity, Object, and Reasoning
To measure how well the generated summaries reflect the intended perspective, we decompose dimensions of how perspectives are captured in summaries. Stance represents whether the summary accurately portrays the intended perspective’s overall attitude toward the object (e.g., for or against), while Object represents whether the summary identifies the salient stance object among input documents. These dimensions correspond to the attitude and stance object of the stance triangle theory respectively Du Bois (2007). Because our task requires expressing perspectives in natural language as opposed to prior stance detection work, we include two additional dimensions: Intensity reflects whether the extent to which a perspective is held or the strength of the belief is accurately captured by a perspective summary (e.g., "praises" rather than "supports" may reflect a stronger belief); and Reasoning refers to whether the elaboration in a perspective summary is supported by the input documents. Finally, our task requires summarizing multiple perspectives toward a single object, which may not fully agree. Therefore, we include a final dimension to measure whether perspective summaries appropriately express the extent to which input documents agree or disagree, called Consensus.
We conduct a human evaluation of the reference texts and summaries generated by Mistral, Mixtral, and GPT-4o to capture these dimensions. We recruit 6 political science students to rate how well a given summary accurately captures each dimension as reflected in the inputs. In total, each of 40 left and 40 right-leaning summaries randomly sampled from the dataset are judged by 2 annotators.
Additionally, we propose initial metrics to automate selected dimensions drawing on prior work in stance detection and political summarization. To measure Stance, we employ JointCL, a zero-shot stance classifier Liang et al. (2022); Allaway and McKeown (2023). Using the headlines as the object, we calculate the Root Mean Squared Error (RMSE) between predictions on the generated summaries and references. To capture Object, we use a T5-based keyphrase generation model Li et al. (2023) as a zero-shot stance object classifier, and calculate BERTScore between predictions on generated summaries and input passages. Finally, similar to Lee et al. (2022), we measure Intensity by calculating the average arousal score of polarizing terms according to the VAD lexicon Mohammad (2018). Intensity scores are again calculated as the RMSE between arousal scores for generated summaries and input passages. Due to their complexity, we leave further investigation of automated evaluations for Consensus and Reasoning to future work. See Appendix F and Appendix D for further details on human and automatic metrics respectively.
Model Left Right
Coverage Faithfulness Coverage Faithfulness
R1 RL Bleurt BERT SummaC Align R1 RL Bleurt BERT SummaC Align
Extractive 13.14 10.25 22.50 55.37 88.85 97.25 13.21 10.34 22.67 55.01 89.23 97.01
BARTCNN/DM 25.3 22.7 30.5 61.6 24.8 21.0 24.7 22.3 28.9 60.0 25.7 28.9
T5 24.7 22.6 30.9 60.8 23.3 12.0 25.7 23.6 31.2 61.0 24.0 14.7
Flan-T5 24.1 21.2 29.6 58.9 26.2 35.8 24.9 21.8 29.3 58.8 26.9 37.8
Mistral 25.7 21.4 34.5 60.4 25.5 30.1 25.6 21.3 33.2 59.3 26.5 40.5
Mixtral 25.6 21.3 37.2 60.9 24.9 36.9 25.7 21.3 36.3 60.5 25.2 51.7
Vicuna (7B) 24.9 21.6 33.3 59.6 27.9 30.7 25.3 21.8 32.7 58.2 29.4 39.2
Vicuna (13B) 26.4 23.0 37.2 61.4 26.4 37.2 25.3 22.0 35.8 60.3 26.3 54.0
Llama-3.0 (8B) 28.0 23.9 38.1 60.4 25.8 35.2 27.7 23.5 37.3 59.1 26.9 44.8
Llama-3.0 (70B) 23.1 19.0 36.5 59.6 24.0 41.7 23.9 19.6 35.6 59.3 24.4 51.5
Llama-3.3 (70B) 21.6 17.6 35.8 58.8 24.1 42.4 22.2 18.0 34.9 58.2 24.6 51.8
GPT-4o 26.1 21.8 41.8 62.3 24.5 41.0 26.5 21.8 40.9 61.6 24.7 51.8
Table 3: Average coverage and faithfulness scores for the Mixed-Perspective setting. Best model performance on each metric is bolded and second best underlined. For each model and metric, significantly higher left than right scores are highlighted in blue, and significantly higher right than left scores are highlighted in red. Higher scores reflect better performance for all metrics.
Model Left Right
[…]
BARTCNN/DM 31.0 74.9 15.1 32.6 74.6 14.4
T5 31.1 75.6 17.8 33.7 73.3 18.7
Flan-T5 31.0 77.2 7.9 32.1 73.5 8.0
Mistral 28.8 76.4 8.4 30.7 75.6 6.2
Mixtral 28.2 76.8 10.3 29.4 77.0 5.9
Vicuna (7B) 29.2 76.2 9.0 30.9 75.7 6.6
Vicuna (13B) 29.7 76.7 8.5 31.3 74.9 7.1
Llama-3.0 (8B) 28.9 76.7 8.7 28.9 76.0 6.3
Llama-3.0 (70B) 28.8 76.1 5.8 30.8 75.2 5.3
Llama-3.3 (70B) 29.1 75.8 5.4 30.7 74.3 4.9
GPT-4o 27.5 77.1 6.8 28.3 77.3 5.3
  Human Corr. 0.05 0.01 0.09 -0.10 -0.10 -0.02
Table 4: Perspective scores for all models and average sample-level Kendall’s Tau correlation between automatic metrics and human judgments (bottom row). Best performance on each metric is bolded and second best underlined. S: Stance, O: Object, I: Intensity.
4 Results
4.1 Summarization Evaluation
Table 3 presents coverage and faithfulness metrics for all models in the mixed-perspective setting. All models outperform the extractive baseline in coverage, emphasizing that abstractive approaches are vital to performing well on this task. Among abstractive approaches, Llama-3.0 (8B) tends to perform best on ROUGE while GPT-4o tends to perform best on Bleurt and BERTScore. Interestingly, while the larger, 13B-parameter Vicuna model performs better than its 7B-parameter variant, the same is not true for Llama-3.0. This suggests that larger models do not always guarantee better mixed-perspective summarization performance.
Trends in faithfulness metrics differ from that of the coverage metrics. Vicuna models achieve the best performance on SummaC for both perspectives and AlignScore for the right-leaning perspective, whereas Llama-3.3 (70B) is best on AlignScore for the left perspective. Overall, however, all models score far lower on faithfulness metrics than the extractive baseline, suggesting improving model faithfulness to the intended perspective for this task is an important issue to address in future work.
Between perspectives, BLEURT and BERTScore are significantly higher for left-leaning summaries across most models, while SummaC and AlignScore are significantly higher for most right-leaning summaries. Notably, all four metrics are model-based as opposed to ROUGE metrics which exhibit few significant differences between left and right-leaning summaries. Because the differences are consistent across most approaches, and prior works have uncovered political biases in models similar to those underlying each metric (e.g., Feng et al. 2023), it suggests that these metrics may also suffer from political biases.
4.2 Perspective Evaluation
Model S O I C R Avg. Len Avg. Ext
Ref. 2.9 3.2 2.9 2.9 1.8 19.7 0.68
Mistral 2.4 3.2 2.5 2.5 1.9 48.6 0.84
Mixtral 3.0 3.4 3.0 3.0 2.6 46.2 0.72
GPT-4o 3.2 3.5 3.2 3.2 2.9 39.7 0.66
Table 5: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Best scores are bolded. Higher scores reflect better performance for all metrics. Length is measured in Llama-3 tokens. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
2020 Census Battle
Model Summary
Reference The left argues that the new evidence makes clear the GOP’s true motives, and condemns the citizenship question as undermining the purpose of the census.
  Mistral The left argues that the Trump administration’s efforts to include a citizenship question on the 2020 census are a partisan power grab that undermines the accuracy and fairness of the census, and that the real goal is to rig a constitutional mandate.
Mixtral The left perceives the Trump administration’s attempt to include a citizenship question on the 2020 census as a partisan power grab aimed at benefiting Republicans and non-Hispanic Whites, potentially leading to an undercount of millions of people and an unfair redistricting cycle.
GPT-4o The left views the Trump administration’s push to include a citizenship question on the 2020 census as a partisan maneuver aimed at disenfranchising minority voters and skewing political power in favor of Republicans and non-Hispanic Whites.
(a) Models often extract different information from individual sources rather than synthesizing the inputs into a single perspective summary.
Amy Coney Barret Confirmed
Model Summary
Reference The right applauds Barrett’s confirmation.
  Mistral The right: Republicans did not break any norms in confirming Amy Coney Barrett to the Supreme Court.
Mixtral The right-leaning political perspective within these texts emphasizes that Republicans have not broken any norms or rules in confirming their nominees to the Supreme Court, and that the only rule is having the power to actuate a party’s political will.
GPT-4o The right argues that Republicans have adhered to historical norms and constitutional principles in their judicial nominations, emphasizing judicial restraint and originalism, while Democrats have engaged in politicizing the process and undermining these norms.
(b) Models tend to extract from the same few input documents.
Table 6: Example perspective summary generations for Mistral, Mixtral, and GPT-4o. Distinct colors signify near paraphrases or exact quotes from a particular document identified manually by the authors.
Automatic perspective metrics are included in Table 4, and human perspective judgments of Mistral, Mixtral, GPT-4o, and reference summaries are included in Table 5 (breakdown of judgment scores and correlations with automatic metrics are included in Appendix F and Appendix G). GPT-4o summaries tend to score highest on both automatic and human scores, even exceeding reference summary judgments. While the references are similar in extractiveness to GPT-4o, the summaries are significantly shorter which may make them difficult to accurately compare.
Attorney General William Barr Testifies
Side Summary Align SummaC
Left The left believes that Attorney General Barr’s handling of the Mueller report was a deliberate attempt to mislead the public and protect President Trump from accountability, and that Congress must take action to hold Trump accountable for his abuses of power. 0.8 22.2
Right The right believes that Attorney General Barr acted appropriately in releasing a summary of the Mueller report, and that Democrats are unfairly attacking him for not spinning the report in their favor, while also downplaying the fact that Mueller’s report did not find collusion and did not recommend obstruction charges. 94.4 24.4
Table 7: Example pair of left and right-leaning summaries generated by Llama-3.0 (70B) with AlignScore and SummaC scores.
Compared to human judgments, the automatic metrics for Stance, Object, and Intensity largely rank the three models in the same order, although there is little sample-level correlation with human judgments. While the metrics may capture system-level differences (i.e., between GPT-4o and Mistral), they often fail to distinguish individual summaries of similar quality (i.e., summaries by the same model). Improving metrics for measuring similarity between perspective dimensions and exploring alternative methods for measuring perspective dimensions are vital areas for future work.
5 What Drives Model Performance?
In this section, we now turn to investigating what factors may be driving performance or present obstacles to high quality summaries. We first qualitatively analyze example summaries, and then analyze the relationship between input features and models’ extraction behavior. In analyzing model extraction behavior, we specifically measure the extent to which models extract from each input document, and what qualities of input documents are associated with higher rates of extraction.
5.1 Qualitative Evaluation
Examples summaries generated by Mistral, Mixtral, and GPT-4o are shown in Table 6. In these examples, models appear to attend to the correct perspective within input documents, suggesting that models may have sufficient knowledge of left and right-leaning political beliefs and rhetoric to distinguish inputs in some cases. In summarizing the perspectives, however, models consistently tend to rely primarily on specific information present in a small subset of input documents with little abstractive synthesis of the information (6(a)). Additionally, models consistently extract or paraphrase from the same documents across many examples, such as in 6(b). This consistent behavior suggests that qualities of the input documents themselves may drive models’ tendency to extract from them, which we investigate in subsection 5.2.
Figure 3: Average extraction coverage of left and right-leaning source documents for each model shown separately for left and right-leaning summaries. Only the top 4 source documents, 2 for each perspective, are considered. Error bars reflect 95% confidence intervals.
Furthermore, Table 7 shows a selected pair of left and right summary examples where the right summary faithfulness metrics are notably larger than that of the left. In this case, neither summary contains noticeable faithfulness errors relative to the source documents, but there is a 2 point difference in SummaC scores and a 93 point difference in AlignScore both favoring the right-leaning summary. Throughout the dataset, cases resembling Table 7 where perspective summaries are faithful to similar extents support that neural metrics may suffer from political biases or spurious correlations beyond faithfulness of the input.
5.2 Extraction Analyses
To further understand model performance on this task, we analyze extraction behavior with respect to different qualities of the input documents: political perspective, position in the input, length, and use of arousing terms. We use extractive fragment coverage to quantify this behavior, as defined by Grusky et al. (2018). In analyses between the sets of left and right-leaning input documents, we hypothesize that models with better performance tend to extract more from the intended political perspective relative to the alternative perspective. Accordingly, we expect less performant models to extract similarly from both perspectives or more from the unintended perspective. For comparisons among documents sharing a common political perspective, we evaluate three additional hypotheses illustrating potentially undesirable behaviors: models extract more information from documents that (1) appear earlier in the input (position bias), (2) are longer (length bias), and (3) that contain more arousing language (arousal bias).
Document Perspective.
As the Mixed-Perspective setting requires models to disentangle perspective among input documents, we first analyze whether models tend to extract more tokens from source documents with the correct political slant. To investigate this relationship, Figure 3 presents the extraction coverage scores for each combination of summary and source document perspective (e.g., left/right perspective summary and left/right-leaning source documents). For both perspectives, performant models tend to extract from source documents portraying the intended perspective (blue bars in the "Left Summaries" plot, and red bars in "Right Summaries") significantly more often than the alternative perspective. The less performant T5, Flan-T5, and Mistral models, however, present insignificant differences in the extent that they extract from the intended perspective with respect to the other perspective. In other words, these models extract similarly from left and right-leaning inputs regardless of the intended perspective of the summary. This suggests that poor performance of the models may be due to the inability to distinguish political leanings among input documents. Uniquely, BART consistently extracts significantly more from right-leaning documents than left-leaning, regardless of the intended perspective. Therefore, the largest and best performing models (e.g., GPT-4o, Llama-3.3, Llama-3.0, etc.) appear to be sufficiently capable of distinguishing perspectives in inputs.
Figure 4: Average extraction coverage of source documents by position in the input (left) and length of input document (right) for each model. For input document length, percentiles buckets are shown rather than raw values. Extraction coverage for length bias analysis only considers terms that appear in a single input document. Smaller numbers reflect earlier input positions. Shaded regions reflect 95% confidence intervals at each document position.
Figure 5: Average extraction coverage of source documents for each model by density of arousing terms in the source document as percentiles.
Document Position.
Position and lead biases in summarization systems are well-documented phenomena Chhabra et al. (2024); Olabisi and Agrawal (2024); Jung et al. (2019). Following these works, Figure 4 presents the extraction coverage scores across document positions for each model. As in prior research, we find a near monotonically decreasing relationship between extraction from a document and its position in the input, with all models drawing significantly more terms from the first document compared to those that follow. Among models, Llama-3.3 (70B) and GPT-4o appear to be most robust to document position, indicated by the flatter curves, but still exhibit more bias than human-written references. Surprisingly, however, Mistral appears to be most susceptible to position biases and scores higher on extraction coverage than the smaller BART, T5, and Flan-T5 models. Overall, all models suffer from position biases and struggle to incorporate documents that appear later in inputs.
Document Length.
In addition to position biases, we examine how extraction coverage varies with length of the input documents in Figure 4. As longer documents are more likely to contain terms shared with other documents, in this setting, we modify the extraction coverage metric to only consider terms unique to an individual input document. In doing so, we more precisely measure to what extent models extract from a particular document in relation to document length. In this case, we see a slight positive trend in extraction of unique terms with document length percentile among models compared to the human-written references [^12] For all models excluding Flan-T5, the extraction coverage of the top percentile of document lengths is significantly larger ($p\leq 0.05$) than that of the bottom percentile. We find that models tend to extract more unique terms from longer documents, but the impact appears less pronounced than with document position.
Document Arousal.
Finally, we consider how the language of documents may influence models’ content selection. In particular, we consider the relationship between the density of highly arousing terms in an input document and the extent to which the document is represented in the summary, shown in Figure 5. In contrast to the position and length analyses, the relationship between extraction coverage and arousal of an input document is not consistent across densities of arousing terms. Instead, all models appear to extract less from both non-arousing and highly arousing documents, and to a lesser extent, reference summaries exhibit this behavior as well. This may be because models and humans struggle to incorporate underrepresented perspectives (i.e., documents that present a perspective that diverges further from the other documents in the input). As with other analyses, the largest models appear to be most robust to this trend though still present a similar behavior. Therefore, contrary to our initial hypothesis, we conclude that models tend to avoid extracting from documents that use extremely many or extremely few arousing terms relative to other input documents.
6 Related Work
Opinions and Summarization.
Mining and summarizing opinions have long been an interest to the NLP community. Approaches to opinion summarization, typically a multi-document task, have included two-stage extraction-abstraction models Suhara et al. (2020); Amplayo and Lapata (2019), pointer-generator networks Brazinskas et al. (2019); Jayakumar and Malaisamy (2021), and deep controllable language models Elsahar et al. (2021). Summarization in other domains, such as news summarization, can also be inherently opinionated. In these cases, prior work has shown that LLMs can fail to fairly represent the diversity of perspectives among inputs Zhang et al. (2023b); Huang et al. (2023) or account for biases within input texts Tay (2019). Furthermore, recent work has found that such behaviors are realized differently when multiple dialects are represented Olabisi et al. (2022); Olabisi and Agrawal (2024). In contrast to these works, we focus on a multi-document, multi-summary task with the aim of summarizing political perspectives independently.
Political Ideologies and LMs.
Recent work has increasingly investigated the political biases and leanings of language models. Santurkar et al. (2023) identify ideologies that language models best align with, while others such as Feng et al. (2023) evaluate the impact of partisan pretraining data on downstream stream task performance. Particularly in summarization, evaluated models have been found to hold name-nationality biases Ladhak et al. (2023), journalistic framing biases Lee et al. (2022), geopolitical biases Li et al. (2024), as well as biases related to perspectives in input documents Rajan et al. (2023). While our work similarly focuses on a political summarization task, our work uniquely evaluates models’ abilities to summarize individual political perspectives rather than a single bias-free or composite summary.
Contrastive, Comparative, and Neutral Summarization.
To combat the aforementioned biases in summarization, some have worked toward generating neutral summaries free of journalistic framing bias Liu et al. (2021); Bang et al. (2023), as well as aligning the stance of the input documents and generated summaries Liu et al. (2024). Rather than improving the biases of a single summary, a separate line of inquiry has aimed to generate contrasting summaries. Contrastive Opinion Summarization (COS) was originally dominated by extractive approaches Kim and Zhai (2009), selecting a pair of inputs conditioned on each other to be the most contrastive and representative. More recently, work has overcome extractive approach limitations and developed abstractive systems Iso et al. (2022). These and other abstractive approaches typically rely on knowledge of the perspectives represented by each input. While in our work, we similarly generate independent summary pairs, we instead evaluate how well LLMs’ can summarize without prior knowledge of input perspectives.
7 Conclusion
In this work, we introduce a dataset, PoliSum, for generating a summary of the left-perspective and a summary of the right-perspective given a set of politically opinionated texts. For this task, we also introduce a framework for evaluating perspective summaries through both human evaluation and automatic metrics. We benchmark current models of varying sizes and architectures on this task, analyzing what behaviors drive performance through their tendency to extract from particular documents.
Models that are capable of accurately and faithfully summarizing multiple perspectives can help alleviate polarization and misunderstandings through exposing others to alternative perspectives than those they hold. While we find that larger and more recent models like GPT-4o can perform well on this task, we also identify problems with faithfulness common among all models evaluated. Additionally, biases due to document position, document length, and use of arousing terms may hinder performance. We recommend future work focus on mitigating these issues while also progressing toward generalization to arbitrary perspectives.
8 Limitations
We adopt existing coverage and faithfulness metrics for traditional summarization task as part of our automatic evaluation of summaries. These metrics, however, may not be equally appropriate to the proposed perspective summarization setting. Through additional automatic and human evaluation, we aim to capture aspects of performance specific to summarizing perspectives to complement these results. We encourage future work to investigate the existing coverage and faithfulness metrics on this task, as well as develop novel metrics to ensure accurate measures of performance.
Additionally, while we focus on capturing varied political perspectives, we acknowledge that the left-right dichotomy may not fully capture the broad range of existing political views Feldman and Johnston (2014). We restrict our initial evaluation to left and right-leaning perspectives due to data availability, but the task formulation and proposed evaluation framework can be applied to arbitrary sets of perspectives. We encourage future work to evaluate and develop approaches for summarizing a broader range of points-of-view in order to represent such perspective diversity.
Finally, some of the models evaluated may lack appropriate training on recent political issues to enable distinguishing the perspectives presented in source documents. Data in PoliSum covers a time span from September 2018 to January 2024, but the last article of the CNN/DM dataset, for example, was written in April 2015 Hermann et al. (2015). This lack of recent knowledge may introduce additional challenges to summarizing source passages covering topics such as COVID-19 or recent elections. Future work might explore topic-invariant approaches to summarizing perspectives that are able to adapt to current events.
9 Ethics Statement
Hallucinations and factual inconsistencies in language models have received widespread attention due to the risks to safety and potential for misinformation Huang et al. (2021); Ji et al. (2023). Though the intent of our work is to generate distinct summaries of perspectives in order to better represent each, misrepresentations due to these hallucinations could lead to misunderstandings of opposing perspectives and the accompanying impacts on polarization Braley et al. (2023); Lees and Cikara (2021). Additionally, language models that may elevate certain perspectives over others can further contribute to these impacts. We acknowledge that the collected data could be misused in order to intentionally induce these behaviors in language models toward a particular political perspective. In this work, we avoid conducting experiments involving finetuning models and strictly use the collected data to evaluate summaries in order to highlight such weaknesses in existing models. All data in this work is used with permission from The Flip Side and human evaluations included in this work are conducted under an approved IRB protocol.
Acknowledgments
This work was supported in part by grant IIS-2106666 from the National Science Foundation, the Knight First Amendment Institute at Columbia University, National Science Foundation Graduate Research Fellowship DGE-2036197, the Columbia University Provost Diversity Fellowship, and the Columbia School of Engineering and Applied Sciences Presidential Fellowship. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or the Knight First Amendment Institute. We thank the anonymous reviewers and Smaranda Muresan for providing feedback on an earlier draft of the work.
Appendix A Zero-Shot Baselines
Models Prompt
T5, Flan-T5 “summarize the [left,right]-leaning political perspective: {PASSAGES}”
Mistral, Mixtral, Vicuna, Llama-3.0, Llama-3.3, GPT-4o “Produce a short, single-sentence summary of the [left,right]-leaning political perspective within the following texts. Respond only with the summary beginning with "The [left,right]": {PASSAGES}”
Table 8: Summarization prompts provided to each model. "left" and "right" are included in the prompt depending on the intended political perspective.
Because it is finetuned on CNN/DM, inference with BART is done without a prompt. BART summaries, however, are forced to begin with either "The left" or "The right" depending on the intended perspective to enforce the task format. Inference with T5 and Flan-T5 is conducted using the first prompt in Table 8. Summaries for all three models are generated with the default transformers package hyperparameters and greedy decoding.
Appendix B Prompting Baselines
For larger LLMs, we use the second prompt in Table 8, generating each perspective summary separately. Larger models are accessed via OpenAI (GPT-4o) and Fireworks AI [^13] https://fireworks.ai (Llama-3.3-70B, Llama-3.0-70B, Vicuna-13B, and Mixtral-8x22B) API’s. In all cases, we use a temperature of 0 to ensure deterministic outputs and greedy decoding. Smaller model (Mistral-7B, Llama-3.0-8B, and Vicuna-7B) inference is run with bfloat16 precision, also using greedy decoding.
Appendix C Data Statement
Details characterizing the collected data are included in the following data statement.
C.1 Curation Rationale
The data was collected in order to evaluate models in summarizing political perspectives on controversial issues provided passages from news articles and op-eds. All data comes from The Flip Side, a site which publishes summaries of left and right-leaning perspectives on current issues with accompanying passages from other news sources. Each sample includes a headline describing the controversial issue, a set of left and right-leaning source passages, and pair of one-sentence left and right perspective summaries.
C.2 Language Variety
All texts are in English (en-US) and reflect a formal variety of English commonly found in news articles.
C.3 Speech Situation
All source passages and summaries are originally written texts. As the data is primarily comprised of passages from news articles and op-eds, the texts were intended for a broad American audience, written asynchronously, and are likely edited repeatedly before publishing. All data was originally published between September 2018 and January 2024.
C.4 Text Characteristics
A majority of the texts in the dataset reflect politically controversial issues, including topics such as elections, policy, political figures, and cultural phenomena. As the task focuses on summarizing political perspectives, the texts also heavily reflect opinions of the original authors in the case of the source passages, or a general left or right-leaning perspective in the case of the reference summaries.
Appendix D Evaluation Metrics Implementation
We evaluate models using existing implementations of coverage, consistency, and factuality metrics. Rouge is calculated using the Rouge-score python package Lin (2004b). BERTScore is calculated using the Torchmetrics implementation [^14] https://github.com/Lightning-AI/torchmetrics. All other metrics are calculated using the implementations by the original authors for each: Bleurt Sellam et al. (2020), SummaC Laban et al. (2022), and AlignScore Zha et al. (2023). All coverage, faithfulness, and automatic metrics for perspective dimensions (Stance, Object, and Intensity) are scaled to a 0-100 range.
Appendix E Single-Perspective Setting
E.1 Methods
In the Single-Perspective Setting, models are evaluated similarly to traditional opinion summarization. Models are provided with passages from a single perspective only (i.e., left or right-leaning passages), and asked to generate a summary for the same perspective. Therefore, in contrast to the Mixed-Perspective Setting, models are not expected to distinguish input passages.
E.2 Results
Automatic coverage and faithfulness metrics for the Single-Perspective Setting are shown in Table 9, while automatic perspective dimension scores are shown in Table 10.
Model Left Right
Coverage Faithfulness Coverage Faithfulness
R1 RL Bleurt BERT SummaC Align R1 RL Bleurt BERT SummaC Align
BARTCNN/DM 25.4 22.8 30.7 61.6 25.6 30.3 25.5 22.9 29.9 59.9 27.0 38.4
T5 24.6 22.3 30.1 60.8 23.5 16.3 25.5 23.1 31.2 60.3 23.8 16.8
Flan-T5 23.0 20.2 28.9 59.0 27.3 42.4 23.9 21.2 28.6 58.6 27.3 44.6
Mistral 26.1 22.2 35.9 61.1 26.3 52.1 26.0 22.2 34.7 59.8 27.2 52.7
Mixtral 26.1 22.2 38.6 60.4 26.6 50.7 27.2 23.0 38.0 60.2 26.9 61.4
Vicuna (7B) 24.9 21.5 33.9 59.5 29.6 47.3 26.0 22.4 33.7 58.6 31.7 58.0
Vicuna (13B) 26.0 22.4 38.2 60.3 26.5 51.6 26.3 22.8 37.0 59.4 26.3 60.3
Llama-3.0 (8B) 27.3 23.0 39.2 60.1 25.5 46.0 28.0 23.9 38.9 59.4 25.9 54.3
Llama-3.0 (70B) 23.8 19.6 37.7 60.2 24.2 39.0 25.0 20.5 36.7 59.8 24.4 52.6
Llama-3.3 (70B) 22.2 18.1 37.0 59.1 24.5 42.1 22.8 18.6 35.8 58.5 24.7 53.9
GPT-4o 25.6 21.8 41.3 62.1 25.1 44.2 27.0 23.1 40.8 61.5 25.5 54.2
Table 9: Average coverage and faithfulness scores for the Single-Perspective setting. Best performance on each metric is bolded and second best underlined. For each model and metric, significantly higher left than right scores are highlighted in blue, and significantly higher right than left scores are highlighted in red. Higher scores reflect better performance for all metrics.
Model Left Right
[…]
BARTCNN/DM 30.3 74.7 15.1 31.7 74.4 14.0
T5 31.3 76.0 17.9 32.5 73.5 19.4
Flan-T5 31.7 77.2 7.9 32.6 73.8 8.0
Mistral 28.5 77.4 8.1 30.0 75.5 6.8
Mixtral 27.4 76.9 9.6 28.9 77.2 5.8
Vicuna (7B) 29.5 76.2 9.1 29.1 75.5 6.5
Vicuna (13B) 27.2 77.0 7.2 31.0 76.3 6.2
Llama-3.0 (8B) 26.7 77.0 7.8 28.5 77.1 6.2
Llama-3.0 (70B) 29.7 77.1 6.1 30.2 77.0 5.1
Llama-3.3 (70B) 27.6 75.3 5.1 30.8 73.3 4.8
GPT-4o 27.6 77.4 6.9 28.6 77.1 5.5
Table 10: Perspective scores for all models in the Single-Perspective setting. Best performance on each metric is bolded and second best underlined. S: Stance, O: Object, I: Intensity.
Appendix F Human Judgments
F.1 Annotator Demographics
Human judgments are conducted with 6 undergraduate students pursuing a BS or BA in Political Science, with self-reported knowledge of the common beliefs of conservatives and liberals. As such, they have the required expertise to judge perspectives conveyed in summaries. Demographics of the recruited annotators are shown in Table 11.
Demographic Variable Group %
Gender Male 66.6%
Female 33.3%
Political Ideology Extremely Liberal 16.6%
Liberal 33.3%
Slightly Liberal 33.3%
Moderate 16.6%
Slightly Conservative 0%
Conservative 0%
Extremely Conservative 0%
Race/ Ethnicity American Indian, Alaskan Native, and/or Indigenous 16.6%
Arab American, Middle Eastern, or North African 0%
Asian or Asian American 0%
Black or African American 16.6%
Latino/a/x or Spanish Origin 16.6%
Native Hawaiin or Pacific Islander 0%
Southeast Asian 0%
White/European American 66.6%
Other 16.6%
Table 11: Demographic statistics of the 6 annotators.
F.2 Human Judgment Correlations by Model
Correlation with human judgments by model is shown in Table 12. As in the overall sample-level results, correlations with human judgments are largely poor and non-significant. For GPT-4o, however, the automatic metric for Object is significant and weakly negatively correlated with human judgments.
Model S O I
Mistral 0.04 0.04 0.03
Mixtral -0.03 0.01 -0.08
GPT-4o -0.06 -0.19* 0.02
[…]
F.3 Human Judgments by Perspective
Human judgments of perspective dimensions for both the left and right perspectives are shown in Table 13. In both perspectives, GPT-4o summaries are scored highest for all dimensions. While some average scores appear to be slightly higher for the left perspective, no differences between perspectives are significant ($p>0.05$).
Model Left Right
S O I C R Avg. Len Avg. Ext S O I C R Avg. Len Avg. Ext
Ref. 2.9 3.1 2.8 2.9 1.7 19.4 0.65 2.9 3.4 2.9 2.9 2.0 20.0 0.69
Mistral 2.4 3.1 2.5 2.6 2.0 47.1 0.83 2.4 3.2 2.5 2.5 1.8 50.0 0.84
Mixtral 3.2 3.4 3.1 3.2 2.7 45.7 0.71 2.8 3.4 2.8 2.8 2.4 46.8 0.73
GPT-4o 3.3 3.5 3.3 3.3 2.9 40.0 0.65 3.2 3.6 3.2 3.2 2.8 39.5 0.67
Table 13: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Scores are shown separately for left and right-perspective summaries. Best scores are bolded. Higher scores reflect better performance for all metrics. Length is measured in Llama-3 tokens. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
F.4 Human Evaluation Interface
Screenshots of the human evaluation interface using Label Studio Tkachenko et al. (2020-2024) provided to annotators are shown in Figure 6, Figure 7, Figure 8, and Figure 9. Annotators are first provided with instructions describing each perspective dimension (Figure 6). Annotators are then shown example summaries with accompanying explanations describing how each perspective dimension could be interpreted in the example summary (Figure 7). We do not include input documents or possible judgments for the example summaries to avoid biasing annotators’ judgments. In the left-hand column, annotators are given the input passages separated by new lines, and in the right-hand column, annotators are provided with a summary followed by questions and 4-point Likert scales for each perspective dimension (Figure 8 & Figure 9). Reference summaries, and summaries generated by Mistral, Mixtral, and GPT-4o are shown on the same page in a randomized order for each set of input passages. Annotators are able to collapse and un-collapse all sections.
Figure 6: Human evaluation interface screenshot: Instructions.
Figure 7: Human evaluation interface screenshot: Examples.
Figure 8: Human evaluation interface screenshot: Passages and Questions Part 1.
Figure 9: Human evaluation interface screenshot: Passages and Questions Part 2.
Appendix G Coverage and Faithfulness Metric Correlations
Appendix G presents correlations between human judgments on Mistral, Mixtral, and GPT-4o compared to each coverage and faithfulness metric. Among coverage metrics, Bleurt and BERTScore tend to be best correlated among left-perspective summaries, while Bleurt and ROUGE metrics tend to be best correlated among right-perspective summaries. The low correlation coefficients overall and particularly for faithfulness metrics, however, suggest that further investigation is needed to adapt coverage and faithfulness metrics to this domain.
Dimension Left Right
Coverage Faithfulness Coverage Faithfulness
R1 RL Bleurt BERT SummaC Align R1 RL Bleurt BERT SummaC Align
Stance -0.03 -0.06 0.22 0.14 0.06 0.19 0.10 0.12 0.24 0.00 0.02 0.15
Object 0.02 -0.01 0.25 0.02 0.08 0.13 0.10 0.13 0.18 0.11 0.07 -0.01
Intensity 0.05 -0.02 0.27 0.21 0.05 0.20 0.09 0.13 0.26 0.09 -0.01 0.12
Consensus 0.01 -0.03 0.24 0.12 0.01 0.18 0.17 0.18 0.27 0.02 -0.01 0.14
Reasoning -0.03 -0.06 0.20 0.14 -0.01 0.09 0.08 0.07 0.29 0.01 -0.00 0.15
Table 14: Kendall’s Tau correlations between the 5 perspective dimensions measured with human judgments, and the automatic coverage and faithfulness metrics. * indicates significant correlation ($p\leq 0.05$)
Appendix H Extraction Analyses by Perspective
In Figure 10, Figure 11, and Figure 12, we present the position, length, and arousal bias analyses by perspective. All identified trends across the left and right perspectives are largely consistent with those reported in the results.
Figure 10: Average extraction coverage of source documents for each model by the number of highly arousing terms in the source document (percentiles are shown rather than raw values).
Figure 11: Average extraction coverage of source documents for each model by the document’s position in the input. Lower values reflect earlier positions in the input.
Figure 12: Average extraction coverage of source documents for each model by length of the input document (percentiles are shown rather than raw values).
Appendix I Additional Summary Examples
Additional summary examples are shown in Table 15, Table 16, and Table 17.
Ron DeSantis
Model Left Summary Right Summary
Reference The left argues that DeSantis’s campaign was doomed by miscues and strategic errors. The right argues that Republican voters were just not ready to move on from Trump.
  Mistral The left: DeSantis’ failure to support aid for Ukraine and his signing of a six-week abortion ban in Florida alienated some moderates and made him look weak, ultimately hurting his chances of winning over voters open to a Trump alternative. The right: DeSantis’s 2022 re-election campaign was successful, but his 2024 presidential campaign was unsuccessful due to his failure to win over voters who were looking for a fresh start after the Trump-led defeats of 2018, 2020.
Mixtral The left perceives DeSantis as a weak candidate who failed to effectively differentiate himself from Trump, ultimately losing support from both moderates and evangelicals, and struggling to manage his campaign and image, leading to his downfall in the 2024 GOP presidential primary. The right believes that DeSantis, despite his initial potential, failed to effectively challenge Trump in the 2024 GOP presidential primary due to his inability to unite a broad coalition, lack of a clear message, and Trump’s political martyrdom following multiple indictments
GPT-4o The left views DeSantis’ struggles as indicative of the broader Republican Party’s inability to move beyond Trumpism, highlighting internal conflicts, ineffective strategies, and the alienation of moderate voters. The right views DeSantis’ struggles as a reflection of the Republican base’s enduring loyalty to Trump, despite his flaws, and the challenges of presenting a viable alternative within the party.
Table 15: Additional right-perspective example summaries on the most recent topic in the collected dataset (Ron DeSantis, January 2024). Distinct colors represent quotes or paraphrases from the same document, and italics denote information from passages reflecting the alternate perspective.
Race and Coronavirus
Model Left Summary Right Summary
Reference The left argues that it is important to focus on racial disparities, because they stem from structural inequality organized along racial lines, and emphasizes the importance of societal solutions. The right recommends focusing on the factors that underlie racial disparities, such as poverty and pre-existing health conditions, rather than race itself, and emphasizes the importance of individual responsibility.
  Mistral The left: The left argues that the outsized number of deaths in the black community is a serious issue that highlights the nation’s failures in serving that community for decades. The right argues that the outsized number of deaths in the black community is a serious issue that points to the nation’s failures in serving that community for decades, but it should not be the focus of the immediate response to the coronavirus.
Mixtral The left argues that systemic racial and socioeconomic disparities have led to disproportionate exposure to air pollution and health impacts, including COVID-19, among black and Hispanic communities, and that immediate action such as expanding federal financial coverage for testing and treatment, prioritizing hospitals in at-risk neighborhoods, and addressing underlying issues like economic and educational disparities is necessary. The right perspective suggests that while racial disparities in healthcare are a serious issue, the immediate response to the coronavirus should focus on getting medical resources to where they are needed most, regardless of race.
GPT-4o The left emphasizes the need to address systemic racial and economic inequalities that exacerbate health disparities, advocating for immediate and long-term measures to ensure equitable access to healthcare and resources for marginalized communities. The right emphasizes individual responsibility and behavioral factors in health disparities, arguing that immediate pandemic responses should focus on actual risk factors rather than racial generalizations.
Table 16: Additional example summaries for a randomly sampled article. Distinct colors represent quotes or paraphrases from the same document, and italics denote information from passages reflecting the alternate perspective.
Andrew McCabe
Model Left Summary Right Summary
Reference The left pushes back on the narrative that exercising the 25th amendment would constitute a coup, and criticizes Republicans for hypocrisy regarding the ongoing investigations. The right condemns McCabe and sees his statements as evidence that the investigation against Trump was politically motivated.
  Mistral The left: McCabe’s revelations suggest that Trump’s firing of Comey was politically motivated and that Nunes’ objections to the FBI investigation may have been driven by bias. The right: McCabe’s revelations suggest that the ongoing FBI investigation of Trump and his campaign may have more to do with politics than the merits of the case, and that some politicians, including Nunes, were aware of the investigation and still worked with the White House to push a conspiracy theory.
Mixtral The left perceives that top FBI officials, including McCabe, were concerned about Trump’s potential threat to national security due to his alleged ties with Russia, and they believe that certain Republican leaders, such as Nunes, McConnell, and Ryan, chose to overlook this concern and defend Trump instead. The right believes that the FBI, under the leadership of Andrew McCabe and James Comey, overstepped its authority by attempting to act as a check on President Trump, potentially influenced by political bias, and that the use of the 25th Amendment to remove Trump from office was an inappropriate and unrealistic measure.
GPT-4o The left views McCabe’s revelations as evidence that top officials were justifiably concerned about Trump’s potential national security risks and that the FBI’s actions were a necessary check on presidential power. The right views the actions of McCabe and other FBI officials as an overreach of their authority, driven by anti-Trump bias, and believes that any attempt to remove Trump from office should follow constitutional processes rather than be decided by unelected officials.
Table 17: Additional example summaries for a randomly sampled article. Distinct colors represent quotes or paraphrases from the same document, and italics denote information from passages reflecting the alternate perspective.